OpenStudy (anonymous):

George measured the capacity of a cardboard box as 400 in3. The true capacity of the cardboard box is 450 in3. What is the percent error in George’s measurement? If necessary, round to the nearest percent. %

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

per cent change =(original-new)/original =400-450 =50/400 *100=1.25% error